OpenStudy (anonymous):

what's the probability that x=4 in the binomial experiment where n=9 and p=.3?

OpenStudy (anonymous):

Consider that \(x=4\) with \(n=9\) implies four out of nine Bernoulli trials ended in success; the number of ways this could occur is given by \(_9C_4=\dfrac{9!}{5!4!}=\frac{9\times8\times7\times6}{4\times3\times2}=9\times2\times7=126\). Given each trial is independent, and the probability for a trial to end in success is \(p=0.3\) and failure \(1-p=0.7\) the probability of any particular way of getting four out of nine trials is \(p^4(1-p)^{9-4}=0.3^40.7^5\) and counting all the different possible ways gives a final probability of \(_9C_4p^4(1-p)^{9-4}=126(0.3)^4(0.7)^5\approx0.172\) i.e. \(17.2\%\).

OpenStudy (anonymous):

In general, for \(X\sim\text{Binom}(n,p)\) we have \(\mathbb{P}(X=x)=\,_nC_xp^x(1-p)^{n-x}\)